OpenStudy (anonymous):

Find the derivative of f(x) = -9/x at x = 12. 3/4 16 4/3 1/16

OpenStudy (anonymous):

okay, this one is easy if you know how to derivate \[x^n\]

OpenStudy (anonymous):

what you have to understand is that saying \[-9\div x\] is the same as\[(-9)1\div x\] and that \[(-9)* 1\div x=(-9)*x ^{-1}\]

OpenStudy (anonymous):

and you know that \[(ax^n)'=nax ^{n-1}\]

OpenStudy (anonymous):

in this case, a=-9 and n=-1

OpenStudy (anonymous):

so, can you tell me the answer?

OpenStudy (anonymous):

Where do I plug it in at?

OpenStudy (anonymous):

see the formula I just posted

OpenStudy (anonymous):

you have to use the one of \[(x^n)'=nx^{n-1}\]

OpenStudy (anonymous):

Okay so I got -12

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

see

OpenStudy (anonymous):

\[(ax^n)'=nax ^{n-1}, -9=a, n=-1, (-9x ^{-1})'=(-1)(-9)x ^{-1-1}=9x ^{-2}\]

OpenStudy (anonymous):

Oh so would it be 1/16?

OpenStudy (anonymous):

hmmm see\[x ^{-2}=\frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

so\[9x ^{-2}=\frac{ 9 }{ x^2 }\]

OpenStudy (anonymous):

if x is 12 then it is 9/144

OpenStudy (anonymous):

so yeah hahaha it's 1/16.... forgot to simplify xD